OpenStudy (anonymous):

Find the exact solutions to each equation for the interval 0<=x<2pi. 2cos^2 x-3 cos x+1=0

myininaya (myininaya):

hey yotis the question before they missed a solution

myininaya (myininaya):

let u=cosx 2u^2-3u+1=0 2u^2-2u-u+1=0 2u(u-1)-(u-1)=0 (u-1)(2u-1)=0 u=1 u=1/2 but remember u=cosx cosx=1 so x=0,2pi cosx=1/2 so x=pi/3, 5pi/3

OpenStudy (anonymous):

so theirs four solutions to this one?

myininaya (myininaya):

we don't include 2pi if x<2pi

myininaya (myininaya):

we include 2p if x<=2pi

OpenStudy (anonymous):

so its just 0, pi/3 and 5pi/3 ?

myininaya (myininaya):

yes! :)

OpenStudy (anonymous):

alright, thanks.

OpenStudy (anonymous):

haha that pic is pretty gay ill admit it but yeah i linked it to my facebook.

myininaya (myininaya):

oh i was going to say u guys look cute (dont worry im not a guy)

OpenStudy (anonymous):

haha thanks

myininaya (myininaya):

lol